03) (2+2+2+1-7 marks) Let X have a Poisson distribution with parameter i-9. a. What value of a such that P(x2 a) 0.005 b. Compute P(3 <x <10) c. Compute P (x9)

plz solv quickly s
Answer #1

Ans:

P(x=k)=e-9*(9k/k!)

a)

P(x>=a)=0.005

P(x<a)=1-0.005=0.995

P(x<=a-1)=0.995

check using excel function or calulator:

poisson(17,9,true)=0.995

P(x<=17)=0.995

a-1=17

a=18

b)

P(3<x<10)=P(x=4)+P(x=5)+P(x=6)+P(x=7)+P(x=8)+P(x=9)

=0.5662

x p(x)
4 0.0337
5 0.0607
6 0.0911
7 0.1171
8 0.1318
9 0.1318
Total 0.5662

c)P(x=9)=e-9*(99/9!)=0.1318